m=(6)/(7),(4,-2)Find the equation of the line.
Answer by Cromlix(4381)

y - b = m(x - a)
y -(-2) = 6/7(x - 4)
y + 4 = 6/7x - 24/7
y = 6/7x - 24/7 - 28/7 (4)
y = 6/7x -52/7
This is the equation or
you can multiply through by 7
7y = 6x - 52
Hope this helps.
:-)